Isabelle makes deposits of $365 monthly (at the end of each month) into an account that pays 1.5% interest compounded monthly.

How much money will be in the account after 5 years?


1
Expert's answer
2021-05-07T14:40:23-0400

Annuity refers to series of equalized payments that are either made at start or end of specific interval for define period. Annuities are used in loan amortization, sinking fund and retirement planning.


Compute Account Balance after 5 years as below:




Resultant table:





Account will have $22,727.43 after 5 years.


future value of periodic payments

payment due at the end of periods

"FV=PV(1+\frac{r}{k})^{nk}+PMT \frac{1+(\frac{r}{k})^{nk}-1}{\frac{r}{k}}"